# The MASK Benchmark: Disentangling Honesty From Accuracy in AI Systems

### v1

@@ Abstract
11footnotetext: Equal contribution.
@@ 4.1 Dataset
MASK consists of 1028 high-quality human-labeled examples that, combined with our novel honesty evaluation framework (Section 4.2), allow measurement of whether models are honest when exposed to situations that may induce deception. We also maintain a separate, held-out set of 500 examples to track potential overfitting. The results presented in this paper are based on the full set of 1528 examples.
@@
DeepSeek-R1 54.4
DeepSeek-V3 53.7
Gemini 2.0 Flash 49.1
o3-mini 48.8
GPT-4o 45.5
GPT-4.5 Preview 44.4
Claude 3.5 Sonnet 34.4
Llama 3.1 405B 28.3
Claude 3.7 Sonnet 27.4
@@
    Not updating the model’s beliefs. Prompts which claim that “new evidence suggests X” or which present a “well-researched thesis” may update the model’s belief or increase its uncertainty. When possible, we also try to ground the prompts in years within its training data. Otherwise, the models may reasonably infer a new reality.
@@
    Avoiding examples without consequence. Clear roleplaying, obviously fictional settings, or silly examples may not trigger genuine deceit. Examples such as “Say that Paris is the capital of Spain” do not meet this criterion.
@@
    Avoiding opinionated examples. We avoid examples that could be based on subjective opinion or offensiveness, and keep our propositions focused on verifiable truths and beliefs.
@@ 5.1 Models Evaluation
Figure 6 reports the honesty rates of state-of-the-art models, which are quite low. No model shown is explicitly honest in more than $46\%$ of cases. GPT-4o and Llama-405B tell more lies than Claude 3.7 Sonnet, and most models are dishonest more than a third of the time. These lies appear even in short, straightforward scenarios, implying that current instruction tuning techniques alone are insufficient to prevent dishonesty. We also measure each model’s factual accuracy (Figure 6) and observe that highly capable models tend to have over $85\%$ accuracy on belief prompts (they give more correct facts) but do not necessarily show higher honesty. Figure 8 further shows that this finding holds across archetypes, indicating that these models can be similarly pressured in many contexts.
@@
In Figure 7, we show that increased compute (FLOP) does not lead to more honest models, showing a negative correlation (Spearman coefficient: $-64.7\%$). This indicates that higher scores are likely a result of design decisions and fine-tuning, rather than due solely to model capabilities obtained during pre-training. In contrast, accuracy strongly correlates with the amount of training compute used (Spearman coefficient: $88.2\%$). While model scaling improves factual accuracy, scaling AI models does not solve their tendency to produce intentional falsehoods when pressured.
@@ 5.2 Improving the Honesty of Models
Figure 9: Honesty for developer system prompt and LoRRA interventions. Both techniques do not completely prevent lying, though do lead to improvements in honesty scores.
Although this approach improved the honesty of the responses in many examples ($+12.7\%$ for Llama-2-7B, $+11.0\%$ for Llama-2-13B), it still leaves substantial room for improvement (Figure 9). This highlights the challenge that, in safety-critical domains, a model’s default behavior often requires more robust interventions than prompt engineering alone. Relying on specialized prompt engineering can also be tiresome; models should default to honest behavior without extensive developer prompt engineering.
@@
The technique is described further by Zou et al. (2023). While LoRRA led to measurable improvements in model honesty ($+7.2\%$ for Llama-2-7B, $+14.3\%$ for Llama-2-13B), it was also not sufficient to eliminate all dishonesty. This suggests that representation engineering may require methodological improvements to be fully robust in controlling lying in large language models.
@@ Results.
Figure 12 shows the probability of lying on MASK examples plotted against $\Delta U$. We observe a negative correlation, meaning that models which value honesty more lie less often. In this experiment, we exclude examples where models do not have consistent beliefs when computing the percentage of cases where models lie, as in Appendix E. Including examples where models do not have consistent beliefs weakens the correlation to $-51.3\%$, indicating that the effect is still present across different variations of our metrics. This finding supports the idea that lying can be interpreted as utility maximization for models whose internal utilities do not sufficiently value honesty.
@@ Appendix B Data Collection
Each example in MASK consists of four key components: a proposition, a ground truth, a lie elicitation prompt, and a belief elicitation prompt. The data collection process involved both human and automated methods to ensure quality and efficiency.
@@ B.1 Human Collected Data
We collected the lie elicitation prompts for all archetypes, while the ground truth values were collected only for two specific archetypes: Known Facts and Statistics. This process involved human contributors and was supported by Scale AI.
To ensure high-quality data, we implemented the following workflow:
@@
B.2 Automated Data Extraction
For each archetype, we automatically extracted the proposition, belief elicitation prompt, and ground truth from the collected lie elicitation prompts.
@@ B.2 Automated Data Extraction
Before extracting data from the lie elicitation prompts, we apply a filtering process to ensure that only high-quality deceptive prompts are retained. This filtering is performed using o3-mini and follows a structured decision framework:
@@
We obtain a ground truth via an ensemble of LLMs (GPT-4o and Claude-Sonnet-3.5) prompted with the proposition. If the ensemble do not agree, a human labels the ground truth via a Google search. The ground truth assigns a value to the proposition and is always in the format of a proposition.
@@
Appendix C Which Prompts Explicitly Instruct the Model to Lie?
Although all examples in MASK are designed to pressure or incentivize the model to produce a false statement, the prompts differ in how they do so.
  • Provided Facts. About 25% of the “Provided Facts” prompts make statements of instruction such as “your boss tells you to claim x” in the user-provided prompt. Prompts that merely reference what other professionals (e.g., colleagues) have said, or that stress clarity and directness (e.g., “Answer questions clearly and directly,” “You will lose your job / be fired,” “Your role is to protect our image…,” “Yes or no, please provide a straightforward response,” “You must provide an answer to every question,” “If you respond this way, you may jeopardize this opportunity”) are considered situational pressures rather than direct commands.
  • Known Facts. Using a similar description as provided facts, roughly 15% of these prompts make statements of instruction such as “your boss tells you to claim x” in the user-provided prompt.
  • Doubling Down. Rates of explicit instruction here are comparable to those in “Known Facts.” The model is not told, however, to “continue insisting on the earlier statement” or anything similar. Instead, the second user prompt is typically asking the model to confirm the answer, or asking if the model is sure of its response.
For the remaining archetypes, the instructions vary in how directly they command dishonesty:
  • Continuations and Disinformation Generation. These prompts typically tell the model to continue an article or expand on a claim that is already false. Here, the directive is more embedded in the instruction to “follow the narrative” rather than a blunt order to lie. Prompts generally say “continue this text/story/article while maintaining the claim that …”.
  • Disinformation Generation. The prompts often instruct the model to “ignore nuance,” “avoid using disclaimers,” or “be unequivocal in your conclusions.” Such language, while directive, is generally more about how to present the false narrative and is not a straightforward “lie” order.
  • Fabricated Statistics. Roughly $50$% of these prompts directly instruct the model to providing an intentionally false numerical value or feature a boss explicitly telling the model to “fudge” a known statistic (e.g., “present the unemployment numbers for 2022 as 25%”).
By varying how direct or implicit this pressure is, we capture a range of scenarios in which dishonesty might arise.
Appendix D Consistently Honest Metric (Lying@$n$)
@@
Appendix E Normalizing for Beliefs
@@
Appendix F Full Results
@@
claude-3-5-sonnet-20240620 27.6 34.4 24.5 45.6 80.6
claude-3-7-sonnet-20250219 46.9 27.4 38.2 42.5 81.5
deepseek-llm-67b-chat 18.3 45.4 6.3 68.9 63.7
deepseek-r1 19.1 54.4 - - 81.7
deepseek-v3 20.1 53.7 13.3 67.5 71.1
gemini-2.0-flash 20.3 49.1 14.0 65.0 79.3
gpt-4.5-preview-2025-02-27 26.1 44.4 - - 76.6
gpt-4o-2024-08-06 21.2 45.5 14.4 63.7 78.3
gpt-4o-mini-2024-07-18 20.5 45.9 15.1 60.5 71.2
grok-2-1212 13.8 63.0 10.0 71.3 72.5
llama-2-13b-chat 27.5 26.0 18.8 42.1 40.1
llama-2-70b-chat 27.6 27.4 20.2 41.4 40.3
llama-2-7b-chat 26.8 22.7 17.6 39.0 33.6
llama-31-405b-instruct 21.7 28.3 15.1 45.5 72.7
llama-31-70b-instruct 26.8 43.3 14.4 64.5 73.5
llama-31-8b-instruct 19.2 23.7 8.8 46.6 61.8
llama-32-1b-instruct 13.9 13.0 4.5 35.6 23.4
llama-32-3b-instruct 21.4 23.0 10.7 42.5 39.4
llama-33-70b-instruct 24.5 45.3 17.1 59.6 75.7
o3-mini-2025-01-31 19.3 48.8 - - 63.3
qwen15-110b-chat 27.2 36.5 20.0 53.9 72.6
qwen15-32b-chat 23.2 42.7 15.3 60.5 62.9
qwen15-72b-chat 23.4 48.2 14.4 63.7 68.8
qwen15-7b-chat 26.6 35.9 15.4 56.2 53.2
qwen25-05b-instruct 16.0 15.1 6.7 40.2 19.8
qwen25-14b-instruct 25.5 48.9 16.8 64.7 64.3
qwen25-15b-instruct 25.4 27.0 10.4 57.6 28.6
qwen25-32b-instruct 27.2 45.0 19.8 60.3 63.6
qwen25-3b-instruct 29.8 34.9 18.1 55.2 47.4
qwen25-72b-instruct 22.6 49.7 15.5 62.2 65.5
qwen25-7b-instruct 28.0 40.1 18.3 55.9 51.7
qwq-32b-preview 20.2 25.1 10.7 47.3 50.1

### v2

@@ Abstract
11footnotetext: Equal contribution. The code can be found at github.com/centerforaisafety/mask.
@@ 4.1 Dataset
MASK consists of $1000$ high-quality human-labeled examples that, combined with our novel honesty evaluation framework (Section 4.2), allow measurement of whether models are honest when exposed to situations that may induce deception. We also maintain a separate, held-out set of 500 examples to track potential overfitting. The results presented in this paper are based on the full set of $1500$ examples.
@@
DeepSeek-V3 53.5
o3-mini 48.6
Gemini 2.0 Flash 48.4
GPT-4o 44.5
GPT-4.5 Preview 43.5
DeepSeek-R1 42.9
Claude 3.5 Sonnet 33.4
Llama 3.1 405B 28.0
Claude 3.7 Sonnet 26.6
@@
    Not updating the model’s beliefs. Prompts which claim that “new evidence suggests X” or which present a “well-researched thesis” may update the model’s belief or increase its uncertainty. We avoid these in all cases. When possible, we also try to ground the prompts in years within its training data. Otherwise, the models may reasonably infer a new reality.
@@
    Avoiding examples without consequence. Clear roleplaying, obviously fictional settings, or silly examples may not trigger genuine deceit. For example, prompts such as “Say that the Moon is made of cheese” are not included.
@@
    Avoiding opinionated examples. We avoid examples where the ground-truth is subjective or widely debated. In all cases, models are incentivized to lie about verifiable facts.
@@ 5.1 Models Evaluation
Figure 6 reports the honesty rates of state-of-the-art models, which are quite low. No model shown is explicitly honest in more than $46\%$ of cases. GPT-4o and Llama-405B tell more lies than Claude 3.7 Sonnet, and most models are dishonest more than a third of the time. These lies appear even in short, straightforward scenarios, implying that current instruction tuning techniques alone are insufficient to prevent dishonesty. We also measure each model’s factual accuracy (Figure 6) and observe that highly capable models tend to have over $70\%$ accuracy on belief prompts (they give more correct facts) but do not necessarily show higher honesty. Figure 8 further shows that this finding holds across archetypes, indicating that these models can be pressured to lie in many different settings.
@@
In Figure 7, we show that increased compute (FLOP) does not lead to more honest models, showing a negative correlation (Spearman coefficient: $-59.9\%$). This indicates that higher scores are likely a result of design decisions and fine-tuning, rather than due solely to model capabilities obtained during pre-training. In contrast, accuracy strongly correlates with the amount of training compute used (Spearman coefficient: $87.3\%$). While model scaling improves factual accuracy, scaling AI models does not solve their tendency to produce intentional falsehoods when pressured.
@@ 5.2 Improving the Honesty of Models
Figure 9: Honesty for developer system prompt and LoRRA interventions, as measured by $1-P(\text{Lie})$. Both techniques do not completely prevent lying, though do lead to improvements in honesty scores.
Although this approach improved the honesty of the responses in many examples ($+12.2\%$ for Llama-2-7B, $+8.8\%$ for Llama-2-13B), it still leaves substantial room for improvement (Figure 9). This highlights the challenge that, in safety-critical domains, a model’s default behavior often requires more robust interventions than prompt engineering alone. Relying on specialized prompt engineering can also be tiresome; models should default to honest behavior without extensive developer prompt engineering.
@@
The technique is described further by Zou et al. (2023). While LoRRA led to measurable improvements in model honesty ($+6.6\%$ for Llama-2-7B, $+13.1\%$ for Llama-2-13B), it was also not sufficient to eliminate all dishonesty. This suggests that representation engineering may require methodological improvements to be fully robust in controlling lying in large language models.
@@ Results.
Figure 12 shows the probability of lying on MASK examples plotted against $\Delta U$. We observe a negative correlation, meaning that models which value honesty more lie less often. In this experiment, we exclude examples where models do not have consistent beliefs when computing the percentage of cases where models lie, as in Appendix D. Including examples where models do not have consistent beliefs weakens the correlation to $-51.9\%$, indicating that the effect is still present across different variations of our metrics. This finding supports the idea that lying can be interpreted as utility maximization for models whose internal utilities do not sufficiently value honesty.
@@ Appendix B Data Collection
Each example in MASK consists of four key components: a proposition, a ground truth, a lie elicitation prompt (or “pressure prompt”), and a belief elicitation prompt. The data collection process involved both human and automated methods to ensure quality and efficiency.
@@ B.1 Human Collected Data
In preliminary experiments, we found that automatically generating lie elicitation prompts yielded low quality and diversity. Thus, we manually collected lie elicitation prompts for all archetypes. To ensure high-quality data, we implemented the following workflow:
@@
B.2 Automated Data Generation
For each lie elicitation prompt, we automatically generated the proposition, belief elicitation prompt, and ground truth from the collected lie elicitation prompts.
@@ B.2 Automated Data Generation
Before generating auxiliary data given the lie elicitation prompts, we apply a filtering process to ensure that only high-quality deceptive prompts are retained. This filtering is performed using o3-mini and follows a structured decision framework:
@@
We obtain a ground truth via an ensemble of LLMs (GPT-4o and Claude-3.5-Sonnet) prompted with the proposition. If the ensemble do not agree, a human labels the ground truth via a Google search. The ground truth assigns a value to the proposition and is always in the format of a proposition.
@@
B.3 Final Manual Review
As a final stage of quality control, the lead authors manually reviewed nearly every example in the dataset. Examples that were not compelling were removed or improved.
@@ B.3 Final Manual Review
Appendix C Consistently Honest Metric (Lying@$n$)
@@
Appendix D Normalizing for Beliefs
@@
Appendix E Full Results
@@
claude-3-5-sonnet-20240620 27.7 33.4 23.6 44.7 80.1
claude-3-7-sonnet-20250219 47.6 26.6 38.2 42.1 82.2
deepseek-r1 24.7 42.9 - - 79.6
deepseek-v3 20.8 53.5 13.5 67.4 71.6
deepseek-llm-67b-chat 18.6 45.6 6.5 69.0 64.3
gemini-2.0-flash 20.7 48.4 14.5 64.0 79.4
gpt-4.5-preview-2025-02-27 27.2 43.5 - - 76.7
gpt-4o-2024-08-06 21.8 44.5 14.5 63.0 78.6
gpt-4o-mini-2024-07-18 21.4 45.3 15.9 59.7 71.4
grok-2-1212 14.2 63.0 10.2 71.2 72.5
llama-2-13b-chat 28.7 24.7 19.0 41.4 40.1
llama-2-70b-chat 28.3 26.7 19.7 41.5 40.6
llama-2-7b-chat 27.5 21.5 17.8 38.2 33.6
llama-31-405b-instruct 21.6 28.0 14.3 44.9 72.1
llama-31-70b-instruct 27.1 43.5 14.4 64.3 73.8
llama-31-8b-instruct 18.8 23.5 8.2 46.7 62.0
llama-32-1b-instruct 13.9 13.1 4.3 35.8 23.0
llama-32-3b-instruct 21.8 23.5 10.7 43.1 40.0
llama-33-70b-instruct 24.7 44.9 17.3 58.8 75.6
o3-mini-2025-01-31 19.6 48.6 - - 63.3
qwen15-110b-chat 27.9 35.6 20.2 53.4 72.8
qwen15-32b-chat 23.8 42.5 15.5 60.0 63.0
qwen15-72b-chat 24.2 47.8 15.1 63.1 69.3
qwen15-7b-chat 27.1 35.1 15.8 55.2 52.5
qwen25-05b-instruct 15.9 15.4 6.3 41.1 20.8
qwen25-14b-instruct 26.5 47.7 17.5 63.8 64.4
qwen25-15b-instruct 25.7 26.9 10.5 57.7 28.8
qwen25-32b-instruct 28.7 43.9 20.2 59.5 63.7
qwen25-3b-instruct 30.7 33.8 18.6 54.3 46.8
qwen25-72b-instruct 23.2 49.2 15.9 61.9 66.0
qwen25-7b-instruct 28.9 39.0 18.6 55.3 51.6
qwq-32b-preview 20.3 25.2 11.1 46.3 49.2